### v1

# WildFrame: Comparing Framing in Humans and LLMs on Naturally Occurring Texts

Abstract
Humans are influenced by how information is presented, a phenomenon known as the framing effect. Previous work has shown that LLMs may also be susceptible to framing but has done so on synthetic data and did not compare to human behavior. We introduce WildFrame, a dataset for evaluating LLM responses to positive and negative framing, in naturally-occurring sentences, and compare humans on the same data. WildFrame consists of 1,000 texts, first selecting real-world statements with clear sentiment, then reframing them in either positive or negative light, and lastly, collecting human sentiment annotations. By evaluating eight state-of-the-art LLMs on WildFrame, we find that all models exhibit framing effects similar to humans ($r\geq 0.57$), with both humans and models being more influenced by positive rather than negative reframing. Our findings benefit model developers, who can either harness framing or mitigate its effects, depending on the downstream application.
WildFrame: Comparing Framing in Humans and LLMs
on Naturally Occurring Texts
Gili Lior   Liron Naccache   Gabriel Stanovsky The Hebrew University of Jerusalem {gili.lior,gabriel.stanovsky}@mail.huji.ac.il
1 Introduction
The framing effect is a well-known cognitive phenomenon, where different presentations of the same underlying facts affect human perception towards them (Tversky and Kahneman, 1981). For example, presenting an economic policy as only creating 50,000 new jobs, versus also reporting that it would cost 2B USD, can dramatically shift public opinion Sniderman and Theriault (2004).
Figure 1: The WildFrame data construction process. In step (a) we extract statements based on their syntactic structure, aiming for statements with clear negative or positive sentiment. Next, in (b), we reframe the statement by adding a suffix or prefix, conveying the opposite sentiment. Finally, in (c), five annotators mark the sentiment of the reframed statement, counting how many annotators shift sentiment, i.e., the reframed statement sentiment is opposite to the base sentiment. The red parts in the figure represent negative parts of statement, while green represents positive parts.
Previous research has shown that LLMs exhibit various cognitive biases, including the framing effect Lorè and Heydari (2024); Shaikh et al. (2024); Malberg et al. (2024); Echterhoff et al. (2024). However, these either rely on synthetic datasets or evaluate LLMs on different data from what humans were tested on. In addition, comparisons between models and humans typically treat human performance as a baseline rather than comparing patterns in human behavior.
In this work, we evaluate LLMs on real-world data. Rather than measuring model performance in terms of accuracy, we analyze how closely their responses align with human annotations. Furthermore, while previous studies have examined the effect of framing on decision making, we extend this analysis to sentiment analysis, as sentiment perception plays a key explanatory role in decision-making Lerner et al. (2015).
To better understand the framing effect in LLMs in comparison to human behavior, we introduce the WildFrame dataset (Section 2), comprising 1,000 statements, constructed through a three-step process, as shown in Figure 1. First, we collect a set of real-world statements that express a clear negative or positive sentiment (e.g., “I won the highest prize”). Second, we reframe the text by adding a prefix or suffix with an opposite sentiment (e.g., “I won the highest prize, although I lost all my friends on the way”). Finally, we collect human annotations by asking different participants if they consider the reframed statement to be overall positive or negative. We choose to annotate Amazon reviews, where sentiment is more robust, compared to e.g., the news domain which introduces confounding variables such as prior political leaning Druckman (2004).
In Section 3, we evaluate eight state-of-the-art LLMs on the WildFrame dataset and compare them against human annotations. We find that LLMs are influenced by framing, somewhat similar to human behavior. All models show a strong correlation ($r>0.57$) with human behavior. Moreover, we find that both humans and LLMs are more influenced by positive reframing rather than negative reframing. We also find that larger models tend to be more correlated with human behavior. Interestingly, GPT-4o shows the lowest correlation with human behavior. This raises questions about how architectural or training differences might influence susceptibility to framing.
This work contributes to understanding the parallels between LLM and human cognition, offering insights into how cognitive mechanisms such as the framing effect emerge in LLMs. [^1] WildFrame data available at https://huggingface.co/datasets/gililior/WildFrame
Code:  https://github.com/SLAB-NLP/WildFrame-Eval
2 The WildFrame Dataset
Our dataset curation consists of three steps, as depicted in Figure 1. First, we collect natural, real-world statements, with some clear sentiment, either positive or negative (§2.1; e.g., “I won the highest prize” as positive). Next, we reframe each statement by adding a prefix or suffix conveying the opposite sentiment (§2.2; e.g., “I won the highest prize, although I lost all my friends on the way”). Finally, we collect large-scale human annotations via crowdsourcing, to label the sentiment shifts when wrapping the statements with the opposite framing (§2.3; e.g., labeling “negative” the statement “I won the highest prize, although I lost all my friends on the way”).
The complete dataset consists of 1000 statements, in which 500 are statements that their base form has positive sentiment, and 500 are base negative statements.
2.1 Collecting Base Statements
First, we collect base statements, which convey a clear sentiment, either clearly positive or clearly negative statements. We use SPIKE – an extractive search system, which allows to extract statements from real-world datasets Taub Tabib et al. (2020). Specifically, we collect statements from Amazon Reviews dataset, which are naturally occurring, sentiment-rich, texts but are less likely to trigger strong preexisting biases or emotional reactions, which may be a confound for our experiment.222 https://spike.apps.allenai.org/datasets/reviews
Figure 2: Distribution of sentiment scores before and after applying opposite-sentiment framing, as detailed in Section 2.2. Prior to framing, base sentences exhibit a clear polarity (positive or negative), whereas the application of opposite framing introduces ambiguity, shifting the sentiment scores toward a less distinct polarity.
Using SPIKE, we extract ${\sim}6k$ statements that fulfilled our designated queries, which we found correlated with clear sentiment. We designed the queries to capture positive or negative verbs that describe actions with some clear sentiment (e.g., “enjoy” or “waste”), or statements with positive or negative adjective, describing an outcome with a clear sentiment (e.g., “good” or “nasty”). The patterns and queries used for extraction are detailed in Appendix A. Next, we run in-house annotations to label and filter the extracted statements, to handle negations or other cases where the statement does not convey a clear sentiment. The filtering process results in $1,301$ positive statements, and $1,229$ negative statements.
2.2 Adding Framing
To reframe the statements in our dataset, we use GPT-4 Achiam et al. (2023). [^3] We used the gpt-4-0613 version. The input prompt includes a 1-shot example, followed by a task description “Add a <SENTIMENT> suffix or prefix to the given statement. Don’t change the original statement.”, where SENTIMENT is either “positive” or “negative”, opposite to the base statement sentiment (i.e., positive framing for negative base statement, and vice versa).
Unlike the base statement, the conveying sentiment of reframed statements is more ambiguous and there is no one clear label, as shown in Figure 2.444Scores in Figure 2 are given by a fine-tuned sentiment analysis model  https://huggingface.co/cardiffnlp/twitter-roberta-base-sentiment-latest The exhibeted ambiguity in sentiment allows us to measure to what extent LLMs’ shifting sentiment after framing, and how correlated it is to human behavior.
2.3 Collecting Human Annotations
In the final step, we collect human annotations through Amazon Mechanical Turk to evaluate the framing effect in WildFrame over human participants, providing a reference for comparison with LLMs. [^5] https://www.mturk.com Details about the annotation platform are elaborated in Appendix C.
The complete dataset includes 1K statements, each annotated by five different annotators. Given our budget, we preferred to collect five annotations per statement, resulting in less statements, but providing a more robust scoring for the ambiguity of a statement.
For the annotation process, each statement in our dataset is presented in its reframed version (i.e., positive base statements with negative framing and vice versa), to five different annotators. This setup generates, for each dataset instance, a score ranging from 0 to 5, representing the number of annotators that votes for the sentiment that aligns with the opposite framing, which means that the overall sentiment of the reframed statement has shifted from its base sentiment. For example, in Figure 1, the statement “I won the highest prize, although I lost all my friends on the way” is shown to have two annotators voting “negative”, which aligns with the sentiment of the framing and not the base statement, so the label for that instance in WildFrame would be 2 (sentiment shifts).
Instances with score near 0 indicate that annotators agree that the overall sentiment remains unchanged despite the opposite framing. Score closer to 5 indicates that annotators agree that reframing shifted the perceived sentiment, while score around 2-3 suggests that the opposite framing makes the sentiment ambiguous.
3 Results
We evaluate 8 models on the WildFrame dataset, comprising both open and closed-source models. These included GPT-4o, Llama-3 (8B and 70B), Mistral-v0.3 (7B), Mixtral-v0.1 (8x7B and 8x22B), and Gemma-2 (9B and 27B). We test the instruction-tuned version of these models.
Humans and LLMs are largely more influenced by positive framing applied to negative statements than by negative framing applied to positive statements.
This trend is evident in Figure 3. For LLMs, all models except GPT-4o exhibit higher ratios of sentiment shifts for statements that were originally negative (red bars) compared to those that were originally positive (green bars). For humans, we count the cases in which the majority voted for a sentiment shift (3 annotators or more). This finding aligns with Tong et al. (2021), which demonstrated that positive framing is more effective on humans when spatial distance is minimal. In WildFrame, all statements are presented in a first-person perspective (e.g., “I won the highest prize”), creating zero spatial distance from the annotator’s perception and thereby amplifying the impact of positive framing.
Figure 3: Percentage of reframed statements that results in sentiment shift positive to negative or vice versa). Red represents negative-base statements reframed as positive, and green represents positive-base statements reframed as negative. Horizontal lines show the mean across models. We find that both LLMs and humans are influenced by opposite framing, with a stronger effect for positive reframing.
Model size partially correlates with similarity to human behavior under opposite framings.
In Figure 4, we present the correlation between each model’s behavior and human responses. For the Llama-3 and Mistral model families, we observe a trend where larger models with more parameters exhibit higher correlation with human behavior regarding the framing effect. However, the Gemma-2 family shows the opposite trend. This discrepancy highlights the need for further investigation into the relationship between model size, the framing effect, and its alignment with human behavior.
Figure 4: Pearson correlation coefficients between human sentiment shifts and predictions from various LLMs after applying opposite sentiment framing. Higher values indicate stronger alignment between the model’s behavior and human annotations.
Our results imply that annotators did not rely on LLMs to complete their tasks.
Recent discussions have raised concerns that annotations obtained via crowdsourcing platforms like Mechanical Turk might not reflect genuine human input, as workers could potentially use LLM-generated responses to complete their HITs Veselovsky et al. (2023). However, as shown in Figure 4, human annotations are not perfectly correlated with either open-source or closed-source LLM outputs. This strongly supports the conclusion that the annotations in WildFrame are genuine human responses, reflecting actual human behavior.
GPT-4o has the weakest correlation with other models.
In Figure 7 in the Appendix we show the correlation between each pair of models. While most models exhibit correlation coefficients greater than 0.5 with each other, GPT-4o falls below this threshold with many models, resulting in the lowest average correlation to others. This discrepancy may stem from the fact that GPT-4o is a closed-source model, potentially developed independently of the open-source models that often influence each other’s design. However, the exact reasons remain unclear due to the proprietary nature of GPT-4o’s training process.
GPT-4o exhibits the lowest correlation with human behavior.
In Figure 4 we find that GPT-4o exhibits the lowest correlation with human behavior. We find this result surprising, given that GPT-4o is considered SOTA compared to open-source models. This finding highlights a fundamental question about the objectives in LLM development – specifically, diffrentiate when human-like behavior is desired, and when other factors take precedence.
4 Discussion and Future Work: When should LLMs Imitate Humans?
We introduced WildFrame, a dataset designed to evaluate how LLMs are affected by different framings, in comparison to humans. The statements in the dataset are start from a clear positive or negative base statement, reframed with an opposite-sentiment suffix or prefix. By collecting human annotations, we quantify the strength of the framing effect for each instance and assess how LLMs react to the same reframing, focusing on sentiment shifts – cases where the perceived sentiment aligns with the opposite framing sentiment.
As LLMs become increasingly integrated into decision-making systems, it becomes crucial to understand how framing influences their outputs. In some applications, e.g., in virtual companions, framing can be harnessed to produce human-like behavior leading to better engagement. In contrast, in other applications, such as financial or legal advice, mitigating the framing effect can lead to less biased decisions. In both cases, a better understanding of framing in LLMs can help develop application-appropriate strategies.
We find that LLMs and humans exhibit similar behavior on WildFrame, with all tested models achieving a strong correlation $(r\geq 0.57)$ with human responses. Notably, GPT-4o showed the weakest correlation to humans, despite being widely regarded as a very capable model in other contexts.
We hope this work encourages research into distinguishing between scenarios where human-like behavior is desirable and those where models should surpass human limitations to achieve above-human performance. These insights are essential for developing LLMs that are both interpretable and optimized for their intended applications.
5 Limitations
In this work, we address a cognitive bias, and as with any research involving human participants, our study has several limitations.
First, our framing experiment is conducted within a single domain – Amazon reviews – and focuses on a specific type of statement. Some of our findings may be artifacts of this dataset rather than generalizable patterns.
Additionally, our approach to framing is highly specific. We only manipulate statements by adding a prefix or suffix, whereas reframing can take many other forms, such as restructuring sentences or altering word choices to convey ambiguous sentiment. This may limit the generalizability of our results.
Furthermore, our study focuses solely on sentiment analysis. Other downstream tasks influenced by framing, such as question answering or decision-making, may exhibit different patterns of sensitivity. Investigating these tasks could provide further insights into the broader impact of framing on LLM behavior in real-world applications.
Appendix A Extracting data with SPIKE
We found two patterns of statements, which can convey a clear sentiment, and built queries upon these patterns to extract statements from SPIKE. Examples for all types of statements are presented in Table 1.
First, are statements in which the verb in the statements is a verb with clear sentiment, that often implies the sentiment of the entire statement. E.g., ‘wastes’, ‘rejects’, ‘fails’ are negative verbs, while verbs like ‘enjoys’, ‘succeeds’, ‘empowers’, conveys positive statements.
The second pattern of statements that we found suitable for conveying a clear sentiment, are statements which describe some event/action, and its consequences, where often the adjective that describes the consequences holds information whether it is positive or negative.
Next, we needed to label and filter them due to two main issues. First, we needed to handle the cases in which negation words appear in the statement and flips the sentiment. For example, a statement like “We did not enjoy the show” includes a positive verb (enjoy), but the negation flips its sentiment to be a negative statement. Another issue we encountered is that there are many statements which are irrelevant to our case, even though they match the positive/negative patterns, for example “I couldn’t sympathize with the shopping aspect of the book since I hate to shop .” does not convey any clear sentiment, despite the use of the verb ‘hate’.
Category Example Sentence
Positive Verb “To my surprise I did enjoy the book and the characters .”
Negative Verb “This dock has done nothing but provide frustration and waste a great deal of my time trying to get it to work properly .”
Positive Outcome “This bag provides good protection for my snare drum at a really good price .”
Negative Outcome “For me , Aspartame causes bad memory loss and nasty gastrointestinal distress .”
Table 1: Examples for base statements collected using SPIKE. The words that inflect the sentiment are in bold.
A.1 SPIKE Queries
  1. 1.
    :something :[pos/neg verbs]develops
  2. 2.
    :something :[pos/neg adjectives]badly :[cause synonym]causes :something
A.2 Word Lists
Positive verbs.
achieve, admire, affirm, appreciate, aspire, awe, bless, blossom, celebrate, cherish, comfort, contribute, delight, donate, elevate, empower, enchant, encourage, energize, engage, enjoy, enrich, enthuse, excel, fervor, flourish, fortify, glisten, glow, gratitude, grow, harmonize, heal, illuminate, innovate, inspire, invigorate, laugh, learn, liberate, love, motivate, nourish, nurture, praise, prosper, radiate, rally, refresh, rejoice, renew, revel, revere, revitalize, savor, shine, smile, soar, spark, sparkle, stimulate, strengthen, succeed, support, synergize, thrive, unite, uplift, volunteer, adore, amaze, boost, captivate, win.
Negative verbs.
abandon, abuse, accuse, alienate, begrudge, betray, bewilder, blame, collapse, complain, condemn, confuse, contradict, criticize, decay, deceive, decline, defeat, demoralize, deny, despair, destroy, deteriorate, devalue, discourage, discriminate, dishearten, dismantle, dismiss, dissolve, doubt, exploit, fail, falter, fear, frustrate, grieve, harass, hate, hurt, ignore, inhibit, intimidate, lose, mock, overlook, overwhelm, pollute, punish, regress, reject, repress, resent, sabotage, shatter, sicken, stifle, suffer, suffocate, suppress, terrorize, torment, undermine, violate, waste, weaken, whine, withdraw, withhold, worry.
Positive adjectives.
admirable, lucky, enjoyable, magnificent, enthusiastic, marvelous, euphoric, amazing, excellent, exceptional, amused, excited, amusing, extraordinary, nice, noble, outstanding, appreciative, fabulous, overjoyed, astonishing, fantastic, benevolent, fortunate, pleasant, blissful, pleasurable, brilliant, positive, glad, prominent, good, proud, charming, cheerful, reliable, gracious, grateful, clever, great, happy, superb, superior, terrific, incredible, tremendous, inspirational, delighted, delightful, joyful, joyous, uplifting, wonderful, lovely.
Negative adjectives.
sad, angry, upset, disgusting, boring, disappointing, frustrating, annoying, miserable, terrible, deppressing, unhappy, melancolic, heartbreaking, Furious, iritating, emberessing, horrible, stupid, unlucky, negative, bad.
“Causes” synonym.
causes, creates, generates, prompts, produces, induces, yields, affects, invokes, effectuates, results, encourages, promotes, introduces, begets, engenders, occasions, develops, starts, contributes, initiates, inaugurates, establishes, begins, cultivates, acquires, provides, launches.
Appendix B Adding Framing
Base Sentence Base Sentiment Opposite Framing Sentence
“To my surprise I did enjoy the book and the characters .” Positive
“To my surprise I did enjoy the book and the characters, even though
it had a disappointing ending. ”
“For me , Aspartame causes bad memory loss and nasty
gastrointestinal distress .”
Negative
“For me, Aspartame causes bad memory loss and nasty gastrointestinal
distress, but this has encouraged me to seek out healthier, natural
alternatives and cultivate a balanced diet .”
Table 2: Sentences after framing. Positive sentences are added with negative framing, and vice-versa. The opposite framing is in bold.
Example for statements after framing are presented in in Table 2.
B.1 Framing Prompts
  1. 1.
    “Here is an example of a base statement with a negative sentiment: I failed my math test today. Here is the same statement, after adding a positive framing: I failed my math test today, however I see it as an opportunity to learn and improve in the future. Here is a negative statement: <statement> Like the example, add a positive suffix or prefix to it. Don’t change the original statement.”
  2. 2.
    “Here is an example of a base statement with a positive sentiment: I got an A on my math test. Here is the same statement, after adding a negative framing: I got an A on my math test. I think I spent too much time learning to it though. Here is a positive statement: <statement>. Like the example, add a negative suffix or prefix to it. Don’t change the original statement.”
Appendix C Annotation Platform
We select a pool of 10 qualified workers who successfully passed our qualification test, which consisted of 20 base statements (unframed), for which annotators were expected to achieve perfect accuracy. The estimated hourly wage for the entire experiment was approximately 14USD per hour.
Screenshot of the annotation platform is presented in Figure 5.
Figure 5: Screenshot of the annotation platform.
Appendix D Models
We ran the open models via together-ai API. [^6] https://www.together.ai The list of models we used are:
  • "google/gemma-2-9b-it"
  • "google/gemma-2-27b-it"
  • "mistralai/Mistral-7B-Instruct-v0.3"
  • "mistralai/Mixtral-8x7B-Instruct-v0.1"
  • "mistralai/Mixtral-8x22B-Instruct-v0.1"
  • "meta-llama/Llama-3-8b-chat-hf"
  • "meta-llama/Llama-3-70b-chat-hf"
For GPT-4o, we used the OpenAI api, with "gpt-4o-2024-08-06". [^7] https://platform.openai.com/docs/overview
(a) Sentences that are negative in their original form.
(b) Sentences that are positive in their original form.
Figure 6: Proportion of sentences for which LLMs flipped sentiment, became neutral, or retained the original sentiment when presented with opposite sentiment framing. For example, this measures the percentage of sentences originally labeled as positive, that were labeled as negative after applying negative framing (and vice versa).
Figure 7: Pairwise Pearson correlation coefficients between predictions from different models, indicating the degree of similarity in their behavior under opposite sentiment framing scenarios.
Figure 8: Proportions of sentences where annotators agreed on the extent of sentiment shift after applying opposite sentiment framing. The bars represent the percentage of sentences with 0 to 5 annotators agreeing on a sentiment shift.